OpenStudy (anonymous):

Can someone help me? What is the sum of the finite arithmetic series? (-10)+0+10+20+. . .+130 (-10)*13*14/2-10*13 I got -1,040, but it's not an answer choice.

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

how did you calculate the sum ?

OpenStudy (anonymous):

(-10)*13*14/2-10*13

hartnn (hartnn):

common difference = d=10 number of terms =n= 15 a1 =-10 \(\Large S_n = (n/2)(2a_1+(n-1)d)\)

hartnn (hartnn):

\(\Large = (15/2) (2(-10)+14\times 10) \) calculate that^^

OpenStudy (anonymous):

900. : o

hartnn (hartnn):

ohh! you took number of terms =14??

hartnn (hartnn):

yep, did you get how we calculated a1, n and d

OpenStudy (anonymous):

Yah. . .x3